For a standard normal random variable, what z-score has (a) probability 0.175 to the right? (3 decimal places) (b) probability 0.950 to the left? (3 decimal places)
Answer #1

Solution :

(a)

P(Z > z) = 0.175

1 - P(Z < z) = 0.175

P(Z < z) = 1 - 0.175 = 0.825

P(Z < 0.935) = 0.82

z score = 0.935

(b)

P(Z < z) = 0.95

P(Z < 1.645) = 0.95

z score = 1.645